OpenStudy (anonymous):

Suppose that f is a function with f (25) = 61 and f (25) = 3. If you have no further information, what is your best estimate of f(26.5)?

OpenStudy (anonymous):

So at 25 f has two different values?

OpenStudy (anonymous):

oh it's f(25) = 61 and f'(25) = 3, ask to find f(26.5)

OpenStudy (anonymous):

Using the value of the derivative at 25 and the value of f at 25, find the equation for the tangent line at x = 25 and use it to approximate f(26.5)

OpenStudy (anonymous):

so 61 = 3*25 + b b = -14 f(26.5) = 3*26.5 -14 = 65.5

OpenStudy (anonymous):

Great job!

OpenStudy (anonymous):

thanks ^_^